OpenStudy (anonymous):

–23 – 7.2 A. –30.2 B. –15.8 C. 15.8 D. 30.2

OpenStudy (anonymous):

give me A Madel

OpenStudy (anonymous):

the answer is a

rebeccaxhawaii (rebeccaxhawaii):

add the two

rebeccaxhawaii (rebeccaxhawaii):

but your answer will be a negative number

OpenStudy (anonymous):

Are you sure? Because my book says "-2.5 - 0.7" Rewrite as addition

OpenStudy (anonymous):

d

OpenStudy (anonymous):

"-2.5 - 0.7. Rewrite as addition"*

OpenStudy (anonymous):

okay guesses

OpenStudy (anonymous):

Hmmmm

OpenStudy (anonymous):

leaning towards 15.8

OpenStudy (anonymous):

why

OpenStudy (anonymous):

-23 – 7.2 A. –30.2 B. –15.8 C. 15.8 D. 30.2

OpenStudy (anonymous):

Be cause of wat meh books says

OpenStudy (anonymous):

okay good thought what does your book says

OpenStudy (anonymous):

"Subtract: -2.5 - 0.7" "Solution rewrite as addition. Add." "-2.5 - 0.7 = -2.5 + (10.7) = -3.2

OpenStudy (anonymous):

"

ILovePuppiesLol (ilovepuppieslol):

lmfao @rebeccaxhawaii

OpenStudy (anonymous):

its -30.2